# A short program that computes in the limit a function $f:\mathbb{N}\to\mathbb{N}$ which eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$

### v31

@@ Abstract
It is known that there exists a limit-computable function $f:\mathbb{N}\to\mathbb{N}$ which is not computable. Every known proof of this fact does not lead to the existence of a short computer program that computes $f$ in the limit. For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k}$, $x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
2020 Mathematics Subject Classification: 03D20.
Key words and phrases: computable function, eventual domination, limit-computable function, non-computable function, single-fold Diophantine representation.
1 The Collatz problem leads to a short program that computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$ of unknown computability
@@
3 A short program in MuPAD that computes $f$ in the limit
@@
There exists a set ${\mathcal{A}}(n)\subseteq{\mathbb{N}}^{n+1}$ such that
@@
f(n)f(n)equals the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
@@
For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq\{x_{i}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. From [12] and Lemma 3 in [11], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. A bit shorter program in MuPAD computes $h$ in the limit.

### v32

# Two undecidable decision problems on an ordered pair of non-negative integers

@@ Abstract
For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$. Since $f$ is not computable, no algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not $\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}~(\forall k\in\{0,\ldots,n\}~(1=x_{k}\Rightarrow 1=y_{k}))\wedge$ $(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}+x_{j}=x_{k}\Rightarrow y_{i}+y_{j}=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow$ $y_{i}\cdot y_{j}=y_{k}))$. Similarly, no algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not $\forall~(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists~(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}~(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
2020 Mathematics Subject Classification: 03D20, 11U05.
Key words and phrases: computable function, eventual domination, limit-computable function, single-fold Diophantine representation, undecidable decision problem.
1 The Collatz problem leads to a short computer program that computes in the limit a function $\gamma:\mathbb{N}\to\{0,1\}$ of unknown computability
@@
3 The first undecidable decision problem on an ordered pair of non-negative integers
@@
For every $n\in\mathbb{N}$, there exists a set ${\mathcal{A}}(n)\subseteq{\mathbb{N}}^{n+1}$ such that
@@
For every $n\in\mathbb{N}$, $f(n)$ equals the smallest $b\in\mathbb{N}$ such that every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,b\}^{n+1}$.
@@
Theorem 3.
No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence: every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,m\}^{n+1}$.
Proof.
Since the function $f$ is not computable, it follows from Observation 2. ∎
4 A short program in MuPAD that computes $f$ in the limit
@@
5 The second undecidable decision problem on an ordered pair of non-negative integers
For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $\mathcal{S}$$\subseteq$ $\{x_{j}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. From [12] and Lemma 3 in [11], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. A bit shorter program in MuPAD computes $h$ in the limit.
Theorem 4.
No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence:
$\forall(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}~\exists(y_{0},\ldots,y_{n})\in\{0,\ldots,m\}^{n+1}$
$(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
Proof.
It holds because the function $h$ is not computable. ∎